OpenStudy (anonymous):

find the coordinates of the midpoint of the segment whose endpoints are H(10,1) and K(4,7)

hartnn (hartnn):

Hi @karina123_ \(\Huge \mathcal{\text{Welcome To OpenStudy}\ddot\smile} \) The co-ordinates of midpoint(x,y) on the line segment with end-points (x1,y1) and (x2,y2) is just the average of co-ordinates, given by : \(\large x=\frac{x_1+x_2}{2}\) \(\large y=\frac{y_1+y_2}{2}\) now,just put the values and find the co-ordinates of midpoint (x,y).

hartnn (hartnn):

Did you try finding the mid-point ?

OpenStudy (anonymous):

yes thanks for the help :)

hartnn (hartnn):

your welcome ^_^ if you want me to verify your answer, tell me what you got :)

OpenStudy (anonymous):

(7,4)

hartnn (hartnn):

you're \(\huge \color{red}{\checkmark \quad \ddot \smile}\)